A store sells bluegrass seed that is worth $6 per pound and ryegrass seed that is worth $3 per pound. How much of each should be mixed to obtain 80 pounds of a blend that is worth $5 per pound?
Answer by TimothyLamb(4379)

x = pounds of blue-grass
y = pounds of rye-grass
---
x + y = 80
6x + 3y = 5*80
---
put the system of linear equations into standard form
---
x + y = 80
6x + 3y = 400
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = pounds of blue-grass = 53.3
y = pounds of rye-grass = 26.7
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---